find the radius and interval of convergence of given series
∞
∑ (x+3)^n/2^n
n
=
0
1
Expert's answer
2020-05-20T18:57:52-0400

Given series is "\sum_{n=0}^{\infin} u_n = \sum_{n=0}^{\infin} \frac{(x+3)^n}{2^n}."

So, "u_n = \frac{(x+3)^n}{2^n}"

Now for convergence of series, "\lim_{n\to \infin} |\frac{u_{n+1}}{u_n}| < 1"

"\implies \lim_{n\to \infin} |\frac{(x+3)^{n+1}/2^{n+1}}{(x+3)^n/ 2^n}| < 1"

"\implies \lim_{n\to \infin} |\frac{x+3}{2}| < 1 \\
\implies |x+3| < 2"

So, radius of convergence of given series is 2.